Can I please get help solving this problem? Thank YouMicro Problem Assume a straight-line, downward sloping inverse demand curve, p = 100 - q and a Marginal Costs = 20. 1. What i
Answer #1

1.

To calculate allocatively efficient price, we have:

P = MC

100 - q = 20

q = 80

Price = 20

Quantity = 80

2.

To calculate profit maximizing price, we need:

Revenue = Pq = (100 - q)q

Equating it to MC, we get:

MR = MC

100 - 2q = 20

q = 40

Price = 100 - 40 = 60

Quantity = 40

3.

Efficiency loss (Deadweight loss) with higher price:

= 0.5(Profit maximizing price - MC)(Difference in quantity) = 0.5(60 - 20)(80 - 40) = 800.

If you are satisfied with the answer, please provide a positive rating. Feel free to comment in case of queries.

Have a nice day ahead!